OpenStudy (anonymous):

The area of a trapezoid is given by the formula A 1/2= (b 1 + b 2)h, where base b 1 is parallel to base b 2 and h is the height. Solve the formula for b 2. Show your work.

OpenStudy (anonymous):

A1/2=(b1+b2)h A1/2=bh1+bh2 A1/2-bh1=bh2 bh2=A1/2-bh1 b2=(A1/2-bh1)/2.....!

OpenStudy (anonymous):

Algebraic manipulation!

OpenStudy (anonymous):

actually its b2=(A1/2-bh1)/h